the equation of a line having the given slope and containing the given point: m = -3; (-2,0)
Answer by ankor@dixie-net.com(22740)

the equation of a line having the given slope and containing the given point: m = -3; (-2,0)
The slope/intercept form: mx + b = y
this problem, m=-3, x=-2, y=0
-3(-2) + b = 0
+6 + b = 0
b = 0 - 6
b = -6
then
y = -3x - 6; is the equation